Acapacitor is to be constructed to have a capacitance of 100µF. The area\ of the plates is 6.0mx0.030m and the relative permittivity of thedielectric is7.0.Find the necessary separation of the plates and the electric fieldbstrength if a potential difference of 12V is applied across the capacitor.


1
Expert's answer
2020-07-27T09:35:46-0400

"C=\frac{\epsilon \epsilon_0S}{d}\to d=\frac{\epsilon \epsilon_0S}{C}=\frac{7\cdot8.85\cdot10^{-12}\cdot(6\cdot0.03)}{100\cdot10^{-6}}=0.1\cdot10^{-6}(m)"


"E=U/d=12/(0.1\cdot10^{-6}=120\cdot10^6(V/m)"